Six years ago, Junk Removal Services (JRS) issued high-yield bonds at par with a maturity of 12 years and a face value of $1,000.

Today, the bonds just paid their semi-annual coupons of $70. Suppose six months before the maturity of the bond (right after the bond made its 23rd coupon payment), the yield-to-maturity (APR, semi-annually compounded) is 20%. If you bought a bond right after it made its 23rd coupon payment and held it until maturity, what would be your return on this investment? PLEASE USE FORMULAS AND NOT EXCEL OR A FINANCIAL CALCULATOR.
Answer #1

YTM = C+(F-P)/n

(F+P)/2

0.2 = 70+(1000-P)/1

(1000+P)/2

0.2 =  1070-P  

500+0.5P

100+0.1P = 1070-P

P = 881.81

So, Return on investment = (1000-881.81)/881.81 *100

= 13.40%